OpenStudy (anonymous):

x + y = 15, y + z = 25, and x + z = 20 What is the average (arithmetic mean) of x,y and z ?

OpenStudy (haseeb96):

x+y+y+z+z+x/6 15+25+20/6 =60/6 10 is the average (arithematic mean)

OpenStudy (anonymous):

Haseeb is right. If you add 15+25+20=60 that will be equal to 2(x+y+z) So 30=(x+y+z) And for 3 numbers the average is (x+y+z)/3

OpenStudy (anonymous):

10 Is the average (arithmetic mean).